OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x - 2y = 9 and -3x + 2y = -9 help!!! :)

OpenStudy (bradely):

Add two equations 0=0 No solution

OpenStudy (anonymous):

i said that, but it was wrong, apparently. these are the options: (1, -3) (3, 0) infinitely many solutions (5, 6) no solution i've even tried infinitely many solutions!

OpenStudy (bradely):

no solution

OpenStudy (usukidoll):

what the........ x.x

OpenStudy (anonymous):

if anyone's there can you also help me with this: 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

"Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution."

OpenStudy (anonymous):

0=0 does not mean no solution, but the opposite.

OpenStudy (anonymous):

what does it mean?

OpenStudy (anonymous):

It means you have infinitely many solutions. If you graph it you'll see you will have coincidental lines.

OpenStudy (anonymous):

oh, thank you! can you help me with this - Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (usukidoll):

I also agree with the many solutions too... if there was no solution you wouldn't have an x or a y value that would satisfy the system

OpenStudy (anonymous):

Sorry elijah, but if you want help from me, you have to show me your attempt first :).

OpenStudy (anonymous):

this is what i have, but my teacher said it was still "not quite there" 3x + 4y = -2 2x - 3y = 10 (2)3x +(2)4y = (2)-2 (3)2x - (3)3y = 3(10) 6x + 8y = -4 6x - 9y = 30 6x + 8y = -4 -6x + 9y = -30 =========== 17y = 26 y = 26/17 = 2 2x - 3y = 10 2x - 3(2) = 10 2x - 6 = 10 2x = 4 x = 2

OpenStudy (anonymous):

i think im missing a few steps? im not too sure what i did wrong there

OpenStudy (usukidoll):



OpenStudy (usukidoll):

26/17 is not equal to 2 mate.

OpenStudy (usukidoll):

your y is whacked... and when that happens you will have an incorrect x ok now I'm going to eat... ttyl

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You should've got (2,-2) not sure why you're subtracting in such a way.

OpenStudy (anonymous):

Multiply your second equation by -3, then add to eliminate.

OpenStudy (anonymous):

ohhh thank you so much! i think i get it now :)

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Good attempt though! :p

OpenStudy (usukidoll):

I see the problem.... 6x + 8y = -4 -6x + 9y = -30 sure the x's do cancel and the y is correct it would be 17, but you have a -4 and -30 which would be -4-30 = -34 17y=-34 y = -2

OpenStudy (usukidoll):

2x - 3y = 10 2x - 3(-2) = 10 2x+6=10 -6 -6 2x=4 x=2 ... there we go... now it's perfecto